OpenStudy (anonymous):

form a third degree polynomial function with real coefficients such that 8+i and 7 are zeros f(x) = ???

OpenStudy (anonymous):

Since coefficients are real, the two complex roots have to be conjugates. So the simplest polynomial is \[(x-7)(x -8 -i)(x-8+i)\]

OpenStudy (anonymous):

The question wants the answer in a X^3 +bx^2+cX+# ?? so im confused.....

OpenStudy (dumbcow):

x = 8 +- i x-8 = +-i (x-8)^2 = -1 (x-8)^2 +1 = 0 x^2-16x +65 = 0 --> (x-7)(x^2-16x+65) Now distribute

OpenStudy (anonymous):

@dumbcow, why -i if the given is positive(+)?

OpenStudy (dumbcow):

because complex zeros always come in pairs

OpenStudy (anonymous):

ahh, yeah. . (+) and (-)

OpenStudy (anonymous):

x^3-23x^2-57x-455? is that right?

OpenStudy (dumbcow):

everything except the "-57x" term

OpenStudy (anonymous):

177x?

OpenStudy (dumbcow):

yep:)

OpenStudy (anonymous):

thanks for the direction much appreciated